KingCooper7436:

HELP please! The factors of 52 and the first three factors of 130 are listed in the table below. Fill in the remaining five factors of 130 and then identify the greatest common factor, or GCF, of 52 and 130. Type the multiples in the increasing order.

jesternmais:

do you know how to find L.C.M?

KingCooper7436:

no can i just get the answer

KingCooper7436:

please

jesternmais:

okay so what we are going to do we are going to find the F.C.M for 52 and 130

cashmoneyAX:

anybody need help

KingCooper7436:

idk what that even is

jesternmais:

52=, 52, 104, 156, 208 130= 130 is not a perfect square.

KingCooper7436:

ok

jesternmais:

so now we know that 130 is not a perfect square

jesternmais:

now*

jesternmais:

The factors of 52 and the first three factors of 130 are listed in the table below. Fill in the remaining five factors of 130 and then identify the greatest common factor, or GCF, of 52 and 130. Type the multiples in the increasing order.

jesternmais:

so since we know we can't have any multiples for 130 we are going to leave it alone for now

jesternmais:

now we are going to find the G.F.C for 130

jesternmais:

Greatest common factor (GCF) of 130 and 134 is 2.

jesternmais:

so your answers are 52=52, 104, 156, 208 130 =2, 130 so as you can see there is no G.F.C OR F.C.M

jesternmais:

you will leave as is

jesternmais:

2 will be the answer actually because 2 can go into 52 and 130

jhonyy9:

52 = 2*26= 2*2*13 130 = 2*65= 2*5*13 GCF = 2*13 = 26 @AZ opinion please ? ty